# references.bib

### v1

@@ 5 Ren–Xu counterexamples
for coefficients $t_{i}$ and $a$ in $\mathbf{F}_{3}.$ Since $R_{m+1}$ is not in $Q_{m+d+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ we have $a\neq 0.$ We let
$P=\frac{t_{0}}{a}M_{r-o}x_{3}^{0}+\frac{t_{1}}{a}M_{r-o-1}x_{3}^{1}+\frac{t_{2}}{a}M_{r-o-2}x_{3}^{2}+\frac{t_{3}}{a}M_{r-o-3}x_{3}^{3}+\frac{t_{4}}{a}M_{r-o-4}x_{3}^{4}+\frac{t_{5}}{a}M_{r-o-5}x_{3}^{5}$
so that $T^{\prime}-PR_{m+1}^{\prime}\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$ Since $\deg(P)=r-o=12d-6j+3\geq 9>7$, by Lemma 5.14 such a symmetric polynomial $P$ is attainable with $P$ having degree at most degree $5$ in $x_{3}.$ Since $T^{\prime}$ also has at most degree $5$ in $x_{3}$ and $R_{m+1}^{\prime}$ has degree $0,$ $(-PR_{m+1}^{\prime}+T^{\prime})$ has at most degree $5$ in $x_{3}$. Letting $U=(-PR_{m+1}^{\prime}+T^{\prime})(x_{1}-x_{2})^{2(m+j-1)+1}$, we have $U$ is in $Q_{m+j}(3,\mathbf{F}_{3})$ with degree $3m+6d$ and since $(-PR_{m+1}^{\prime})(x_{1}-x_{2})^{2(m+j-1)+1}$ is generated by $R_{m+1}$ and $T$ is not, $U$ is not generated by $R_{m+1}$. Finally we also have $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ is in $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $33m+6(j-d)+$ 3 . Thus what is left is to prove is $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ and $(-PR_{m+1}^{\prime}+T^{\prime})(x_{1}-x_{2})^{2(m+j-1)+1}$ are generators for $Q_{m+j}(3,\mathbf{F}_{3}).$
Assume for sake of contradiction that $U$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ are not both generators. If $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ is a generator, then any other generator must be of at least degree $3m+6d$ by Lemma 4.3. Yet $U$ is not generated by $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ since it is not generated by $R_{m+1}.$ Thus $U$ must be a generator.
Next, we consider if $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ is not a generator. For $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ to not be a generator there must be a generator in a degree less than $33m+6(j-d)+$ 3. Let it be $G$, and by Lemma 4.3, any other generator must have degree greater than $3m+6d.$ Thus $U$ is not a generator, so $U$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ are both generated by $G$ and specifically $U=QG$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}=SG$ for symmetric polynomials $P$ and $Q.$ Moreover, $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ is the lowest degree generator for $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ so $G=CR_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ for a symmetric polynomial $C.$ This implies $C|\prod(x_{i}-x_{j})^{2},$ and $G$ is not a scalar multiple of $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)},$ so $C$ is a constant. We then have $U$ is a constant multiple of $QR_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)},$ so $U$ is generated by $R_{m+1}$ which is a contradiction.
Thus $U$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ are each generators and together with $1$ they freely generate $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ by Lemma 4.4
@@
Let $Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ have generators $K$ in degree $33m-$ 3 and $T$ in degree $3m$ such that $K$ is not in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$ If $m$ is even then $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by a generator in degree $3m+1,$ $3m+2,$ and $1$
@@
Suppose for the sake of contradiction that $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m.$ Without loss of generality let that generator be $T.$ From [felder2001action], we can let $L^{\prime}$ be a degree $13m+$ 1 generator of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ with coprime integer coefficients. Then $\pi(L^{\prime})\in Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ so $\pi(L^{\prime})$ must be generated by $T$ since any other generator in degree less than degree $13m+$ 1 would violate Lemma 4.3. Moreover, the only degree $1$ symmetric polynomials are constant multiples of $x_{1}+x_{2}+x_{3},$ so we can assume without loss of generality that
@@
Note that from [wang2022explicit] all generators of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ must lie in $\mathbf{Q}[x_{1}-x_{3},x_{2}-x_{3}].$ Thus $(x_{1}+x_{2}+x_{3})T\in\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3}]$ and so $T\in\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3}].$
We also have $T=(x_{1}-x_{2})^{2m+1}T^{\prime}$ for some $s_{12}$-invariant polynomial $T^{\prime}$. Thus by the fundamental theorem of symmetric polynomials $T^{\prime}\in\mathbf{F}_{3}[(x_{1}-x_{3})(x_{2}-x_{3}),x_{1}+x_{2}+x_{3}].$ Note that $\deg T^{\prime}=3m-2m-1=m-1$ and $m$ is even, so $T^{\prime}$ has an odd degree. However, since it is generated by $(x_{1}-x_{3})(x_{2}-x_{3})$ and $x_{1}+x_{2}+x_{3},$ we must have $(x_{1}+x_{2}+x_{3})|T^{\prime}.$ This gives a contradiction because $T$ is a generator.
@@
Assume the claim is true when $m<j$ for some $j\in\mathbf{N}.$ Consider the space $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Let $t$ be the largest natural number less than $j$ such that $t\not\in X$. By the inductive hypothesis $Q_{t}(3,\mathbf{F}_{3})$ has a generator in degree $13t+$ 1 and $23t+$ 2 . By Lemma 5.10, we may let the generators be
$\left((x_{1}+x_{2}+x_{3})\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)-x_{3}B\right)(x_{1}-x_{2})^{2m+1}$
and
$B(x_{1}-x_{2})^{2m+1}$
where $(x_{1}-x_{2})^{2m+1}(x_{1}+x_{2}-2x_{3})A^{\prime}$ and $(x_{1}-x_{2})^{2m+1}B^{\prime}$ are generators for $Q_{t}(3,\mathbf{Q})_{\mathrm{std}}$ and $\pi(A^{\prime})=\pi(B^{\prime})=B.$ From Lemma 5.15, $Q_{t+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by a generator in degree $63t+$ 6 and $33t+$ 3 . Moreover, $R_{t+1}$ is the $33t+$ 3 degree generator by Lemma 5.8. Let $L$ be the degree $63t+$ 6 generator. Suppose $R_{t+1}$ lies in $Q_{t+d}(3,\mathbf{F}_{3}),$ but not $Q_{t+d+1}(3,\mathbf{F}_{3})$ where $d$ is a natural number.
First we consider when $t+d\geq j\geq t+1.$ By Lemma 5.16, $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has generators $R_{t+1}$ and $L\prod(x_{i}-x_{j})^{2(j-t-1)}.$ Note that $R_{t+1}=R_{j}$ by Lemma 5.12, and further $\deg(L)+\deg(R_{t+1})=(6(j-t-1)+3t+6)+3t+3=6j+3.$ By Lemma 4.4, we then have that $R_{t+1}$ and $L\prod(x_{i}-x_{j})^{2(j-t-1)}$ generate $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$
Next, we consider the case where $t+2d-1\geq j\geq t+d+1.$ Notice that by our construction in Lemma 5.15, we can choose $L$ such that it has at most degree $5$ in $x_{3}.$ Thus we can apply Lemma 5.17, which gives us that $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by $R_{t+1}\prod(x_{i}-x_{j})^{2(j-t-d)}$ and a generator in degree $3t+6d.$ Note that $R_{t+1}\prod(x_{i}-x_{j})^{2(j-t-d)}$ is a constant multiple of $R_{j}$ by Lemma 5.12. Moreover, the sum of their degrees is $33t+6(j-t-d)+3+3t+6d=6j+$ 3 as desired.
Finally we consider if $j=t+2d.$ Note that by Lemma 5.17, $Q_{t+2d-1}(3,\mathbf{F}_{3})$ has a generator in degree $3t+6d$ and $33t+6(d-1)+$ 3 . The degree $33t+6(d-1)+$ 3 generator is $R_{t+1}\prod(x_{i}-x_{j})^{2(d-1)},$ and $R_{t+1}$ is divisible by $(x_{1}-x_{2})^{2(t+d)+1}$ where $d$ is maximal, so $R_{t+1}\prod(x_{i}-x_{j})^{2(d-1)}$ does not lie in $Q_{t+2d}(3,\mathbf{F}_{3}).$ Moreover, $Q_{t}(3,\mathbf{F}_{3})$ is a non Ren–Xu counterexample, so $t$ must be even by Lemma 5.6. Then $t+2d$ is even as well, so by Lemma 5.18, $Q_{t+2d}(3,\mathbf{F}_{3})$ has a generator in degree $13(t+2d)+$ 1 and $23(t+2d)+$ 2 .
Now we claim we have exhausted all cases. If we had $j>t+2d,$ since we just showed $t+2d\not\in X,$ we would not have chosen $t$ to be the largest natural number less than $j$ not in $X.$
@@
We can compute the degrees of generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ explicitly. If $m$ has no digits $1$ in its base $3$ representation, then the generators have degree $13m+$ 1 and $23m+$ 2 . Otherwise the lower degree generator is $R_{m}.$ We can deduce the minimal degree of the Ren–Xu counterexamples in $Q_{m}(3,\mathbf{F}_{3}):$ Let $a$ be the greatest natural number such that the $a$th term from the right in the base $3$ representation of $m$ is $1$ Then if $\left\lceil\frac{\left\lceil\frac{m}{3^{a}}\right\rceil-1}{2}\right\rceil=k,$ a minimal degree Ren–Xu counterexample is $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $b=\max\left\{\frac{2m+1-3^{a}(2k+1)}{2},0\right\}.$ The degrees of the generators are then $3^{a}(2k+1)+6b$ and $6m+3-3^{a}(2k+1)-6b$.
6 Representations of $S_{3}$ in $Q_{m}(3,\mathbf{F}_{3})$

### v2

@@ 1. In $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$, for $j\in Z_{\geq 0}$,
for coefficients $t_{j}$ and $a$ in $\mathbf{F}_{3}$. Since $R_{m+1}$ is not in $Q_{m+d+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, we have $a\neq 0$. We let
$P=\frac{t_{0}}{a}M_{r-o}x_{3}^{0}+\frac{t_{1}}{a}M_{r-o-1}x_{3}^{1}+\frac{t_{2}}{a}M_{r-o-2}x_{3}^{2}+\frac{t_{3}}{a}M_{r-o-3}x_{3}^{3}$
$\hphantom{P=}{}+\frac{t_{4}}{a}M_{r-o-4}x_{3}^{4}+\frac{t_{5}}{a}M_{r-o-5}x_{3}^{5},$
so that $T^{\prime}-PR_{m+1}^{\prime}\equiv 0\pmod{(x_{1}-x_{2})^{2}}$ by Lemma 5.13. Since $\deg(P)=r-o=12d-6k+3\geq 9>7$, by Lemma 5.15 such a symmetric polynomial $P$ is attainable with $P$ having degree at most degree $5$ in $x_{3}$. Since $T^{\prime}$ also has at most degree $5$ in $x_{3}$ and $R_{m+1}^{\prime}$ has degree 00, $(-PR_{m+1}^{\prime}+T^{\prime})$ has at most degree $5$ in $x_{3}$. Letting $U=\bigl{(}-PR_{m+1}^{\prime}+T^{\prime}\bigr{)}(x_{1}-x_{2})^{2(m+k-1)+1}$, we have $U$ is in $Q_{m+k}(3,\mathbf{F}_{3})$ with degree $3m+6d$ and since $\bigl{(}-PR_{m+1}^{\prime}\bigr{)}(x_{1}-x_{2})^{2(m+k-1)+1}$ is generated by $R_{m+1}$ and $T$ is not, $U$ is not generated by $R_{m+1}$. Finally, we also have $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ is in $Q_{m+k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $33m+6(k-d)+$ 3. Thus what is left is to prove is $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ and $\bigl{(}-PR_{m+1}^{\prime}+T^{\prime}\bigr{)}(x_{1}-x_{2})^{2(m+k-1)+1}$ are generators for $Q_{m+k}(3,\mathbf{F}_{3})$.
Assume for sake of contradiction that $U$ and $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ are not both generators. If $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ is a generator, then any other generator must be of at least degree $3m+6d$ by Lemma 4.3. Yet $U$ is not generated by $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ since it is not generated by $R_{m+1}$. Thus $U$ must be a generator.
Next, we consider if $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ is not a generator. For $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ to not be a generator there must be a generator in a degree less than $33m+6(k-d)+$ 3. Let it be $G$, and by Lemma 4.3, any other generator must have degree greater than $3m+6d$. Thus $U$ is not a generator, so $U$ and $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ are both generated by $G$ and specifically $U=QG$ and $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}=SG$ for symmetric polynomials $P$ and $Q$. Moreover, $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}$ is the lowest degree generator for $Q_{m+k-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, so $G=CR_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}$ for a symmetric polynomial $C$. This implies $C|\prod(x_{i_{1}}-x_{i_{2}})^{2}$, and $G$ is not a scalar multiple of $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$, so $C$ is a constant. We then have $U$ is a constant multiple of $QR_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d-1)}$, so $U$ is generated by $R_{m+1}$ which is a contradiction.
Thus $U$ and $R_{m+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(k-d)}$ are each generators and together with $1$ they freely generate $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ by Lemma 4.4. ∎
@@
Let $Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ have generators $K$ in degree $33m-$ 3 and $T$ in degree $3m$ such that $K$ is not in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. If $m$ is even, then $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by a generator in degree $13m+$ 1, $23m+$ 2, and $1$.
@@
Suppose for the sake of contradiction that $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator $U$ in degree $13m-$ 1 or $23m-$ 2. Then since $U$ is also in the $-1$ $s_{12}$ eigenspace of $Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, $U$ must be generated by $K$ over $\mathbf{F}[x_{1},x_{2},x_{3}]^{S_{3}}$. Yet $K$ being divisible by a symmetric polynomial violates Corollary 4.2.
Suppose for the sake of contradiction that $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m$. Without loss of generality let that generator be $T$. From [9], we can let $L^{\prime}$ be a degree $13m+$ 1 generator of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ with coprime integer coefficients. Then $\pi(L^{\prime})\in Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, so $\pi(L^{\prime})$ must be generated by $T$ since any other generator in degree less than degree $13m+$ 1 would violate Lemma 4.3. Moreover, the only degree $1$ symmetric polynomials are constant multiples of $x_{1}+x_{2}+x_{3}$, so we can assume without loss of generality that
@@
Note that from [13] all generators of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ must lie in $\mathbf{Q}[x_{1}-x_{3},x_{2}-x_{3}]$. Thus $(x_{1}+x_{2}+x_{3})T\in\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3}]$ and so $T\in\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3}]$.
We also have $T=(x_{1}-x_{2})^{2m+1}T^{\prime}$ for some $s_{12}$-invariant polynomial $T^{\prime}$. Thus by the fundamental theorem of symmetric polynomials $T^{\prime}\in\mathbf{F}_{3}[(x_{1}-x_{3})(x_{2}-x_{3}),x_{1}+x_{2}+x_{3}]$. Note that $\deg T^{\prime}=3m-2m-1=m-1$ and $m$ is even, so $T^{\prime}$ has an odd degree. However, since it is generated by $(x_{1}-x_{3})(x_{2}-x_{3})$ and $x_{1}+x_{2}+x_{3}$, we must have $(x_{1}+x_{2}+x_{3})|T^{\prime}$. This gives a contradiction because $T$ is a generator. ∎
@@
Assume the claim is true when $m<j$ for some $j\in\mathbf{N}$. Consider the space $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Let $t$ be the largest natural number less than $j$ such that $t\not\in X$. By the inductive hypothesis, $Q_{t}(3,\mathbf{F}_{3})$ has a generator in degree $13t+$ 1 and $23t+$ 2. By Lemma 5.10, we may let the generators be
$\biggl{(}x_{1}+x_{2}+x_{3})\pi\biggl{(}\frac{A^{\prime}-B^{\prime}}{3}\biggr{)}-x_{3}B\biggr{)}(x_{1}-x_{2})^{2t+1}\qquad\text{and}\qquad B(x_{1}-x_{2})^{2t+1},$
where $(x_{1}-x_{2})^{2t+1}(x_{1}+x_{2}-2x_{3})A^{\prime}$ and $(x_{1}-x_{2})^{2t+1}B^{\prime}$ are generators for $Q_{t}(3,\mathbf{Q})_{\mathrm{std}}$ and $\pi(A^{\prime})=\pi(B^{\prime})=B$. From Lemma 5.16, $Q_{t+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by a generator in degree ${3t+6}$ and $33t+$ 3. Moreover, $R_{t+1}$ is the $33t+$ 3 degree generator by Lemma 5.8. Let $L$ be the degree $63t+$ 6 generator. Suppose $R_{t+1}$ lies in $Q_{t+d}(3,\mathbf{F}_{3})$, but not $Q_{t+d+1}(3,\mathbf{F}_{3})$, where $d$ is a natural number.
First, we consider when $t+d\geq j\geq t+1$. By Lemma 5.17, $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has generators $R_{t+1}$ and $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-t-1)}$. Note that $R_{t+1}=R_{j}$ by Lemma 5.12, and further $\deg\big{(}L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-t-1)}\big{)}+\deg(R_{t+1})=(6(j-t-1)+3t+6)+3t+3=6j+3$. By Lemma 4.4, we then have that $R_{t+1}$ and $L\prod(x_{i_{1}}-x_{i_{2}})^{2(j-t-1)}$ generate $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$.
Next, we consider the case where $t+2d-1\geq j\geq t+d+1$. Notice that by our construction in Lemma 5.16, we can choose $L$ such that it has at most degree $5$ in $x_{3}$. Thus we can apply Lemma 5.18, which gives us that $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by $R_{t+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(j-t-d)}$ and a generator in degree $3t+6d$. Note that $R_{t+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(j-t-d)}$ is a constant multiple of $R_{j}$ by Lemma 5.12. Moreover, the sum of their degrees is $33t+6(j-t-d)+3+3t+6d=6j+$ 3 as desired.
Finally, we consider if $j=t+2d$. Note that by Lemma 5.18, $Q_{t+2d-1}(3,\mathbf{F}_{3})$ has a generator in degree $3t+6d$ and $33t+6(d-1)+$ 3. The degree $33t+6(d-1)+$ 3 generator is $R_{t+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(d-1)}$, and $R_{t+1}$ is divisible by $(x_{1}-x_{2})^{2(t+d)+1}$, where $d$ is maximal, so $R_{t+1}\prod(x_{i_{1}}-x_{i_{2}})^{2(d-1)}$ does not lie in $Q_{t+2d}(3,\mathbf{F}_{3})$. Moreover, $Q_{t}(3,\mathbf{F}_{3})$ is a non Ren–Xu counterexample, so $t$ must be even by Lemma 5.6. Then $t+2d$ is even as well, so by Lemma 5.19, $Q_{t+2d}(3,\mathbf{F}_{3})$ has a generator in degree $13(t+2d)+$ 1 and $23(t+2d)+$ 2.
Now we claim we have exhausted all cases. If we had $j>t+2d$, since we just showed ${t+2d\not\in X}$, we would not have chosen $t$ to be the largest natural number less than $j$ not in $X$. ∎
@@
We can compute the degrees of generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ explicitly. If $m$ has no digits $1$ in its base $3$ representation, then the generators have degree $13m+$ 1 and $23m+$ 2. Otherwise, the lower degree generator is $R_{m}$. We can deduce the minimal degree of the Ren–Xu counterexamples in $Q_{m}(3,\mathbf{F}_{3})$: Let $a$ be the greatest natural number such that the $a$-th term from the right in the base $3$ representation of $m$ is $1$. Then if $\bigl{\lceil}\frac{\lceil\frac{m}{3^{a}}\rceil-1}{2}\bigr{\rceil}=k$, a minimal degree Ren–Xu counterexample is $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$, where
$b=\max\left\{\frac{2m+1-3^{a}(2k+1)}{2},0\right\}.$
The degrees of the generators are then $3^{a}(2k+1)+6b$ and $6m+3-3^{a}(2k+1)-6b$.
6 Representations of $\boldsymbol{S_{3}}$ in $\boldsymbol{Q_{m}(3,\mathbf{F}_{3})}$